Donald E. Knuth (B.S. and M.S., Case Institute of Technology 1960; Ph.D., California Institute of Technology 1963) is Professor Emeritus of The Art of Computer Programming at Stanford University, where he supervised the Ph.D. dissertations of 28 students since becoming a professor in 1968. He is the author of numerous books, including five volumes (so far) of The Art of Computer Programming, five volumes of Computers & Typesetting, nine volumes of collected papers, and a non-technical book entitled 3:16 Bible Texts Illuminated. His software systems TEX and are extensively used for book publishing throughout the world. He is a member of the American Academy of Arts and Sciences, the National Academy of Sciences, the National Academy of Engineering, and the American Philosophical Society, and he is a foreign associate of the French, Norwegian, Bavarian, and Russian science academies as well as the Royal Society of London. He received the Turing Award from the Association for Computing Machinery in 1974; the National Medal of Science from President Carter in 1979; the Steele Prize from the American Mathematical Society in 1986; the Adelsköld Medal from the Royal Swedish Academy of Sciences in 1994; the Harvey Prize from the Technion of Israel in 1995; the John von Neumann Medal from the Institute of Electrical and Electronic Engineers in 1995; the Kyoto Prize from the Inamori Foundation in 1996; the Frontiers of Knowledge award from the BBVA Foundation in 2010; and the Faraday Medal from the IET in 2011. He holds honorary doctorates from Oxford University, the University of Paris, the Royal Institute of Technology in Stockholm, the University of St. Petersburg, the University of Marne-la-Vallée, Masaryk University, St. Andrews University, Athens University of Economics and Business, the University of Macedonia in Thessaloniki, the University of Tübingen, the University of Oslo, the University of Antwerp, the Swiss Federal Institute of Technology in Zürich, the University of Bordeaux, the University of Glasgow, and nineteen colleges and universities in America.